# Auriga Streams I: disrupting satellites surrounding Milky Way-mass haloes at multiple resolutions

### v1

@@ Abstract
pubyear: 2024pagerange: Auriga Streams I: disrupting satellites surrounding Milky Way-mass haloes at multiple resolutionsB
@@ 1 Introduction
In both the real Universe and the prevailing cosmological paradigm – cold dark matter with a cosmological constant ($\Lambda$CDM) – structure forms hierarchically (Cole et al., 2000) resulting in galaxies surrounded by stellar haloes formed by the remnants of smaller accretions (Searle & Zinn, 1978). As smaller galaxies fall into and orbit around a more massive host, they begin to unravel and form extended stellar streams that eventually phase-mix into a spatially smooth halo (Bullock & Johnston, 2005; Cooper et al., 2010; Gómez et al., 2013).
@@
Figure 1: The accreted stellar mass of the six Auriga galaxies (Au-6, Au-16, Au-21, Au-23, Au-24, Au-27 from left to right) at level 3 resolution, with accretion events split into intact satellites (top), stellar streams (middle), and phase-mixed (bottom) morphologies. The colour corresponds to the amount of stellar mass in a pixel, logarithmically scaled to improve contrast in low/high-density regions. Note that the visible presence of tidal tails from some intact satellites, progenitors for many streams, and cold structure within phase-mixed debris reflects the continuous nature of tidal disruption. The Auriga sample captures a wide range of possible accretion histories for Milky Way-mass haloes, which are efficiently classified by our automated procedure (Section 3.2).
@@
This saga of nearly continuous accretion and disruption of satellites has been observed around both the Milky Way and external galaxies. Wide-area digital sky surveys – in particular the Sloan Digital Sky Survey (York et al., 2000), Dark Energy Survey (DES Collaboration, 2016), and community DECam programs (Drlica-Wagner et al., 2021) – have been incredibly prolific at revealing ever fainter Milky Way satellites (e.g. Willman et al., 2005; Belokurov et al., 2006b; Bechtol et al., 2015; Koposov et al., 2015; Drlica-Wagner et al., 2020; Smith et al., 2024) and stellar streams (e.g. Belokurov et al., 2006a; Koposov et al., 2014; Bernard et al., 2016; Shipp et al., 2018). All-sky astrometry from the Gaia satellite, often in combination with line-of-sight velocities and chemistry from spectrosopic surveys, has produced and characterised an abundance of nearby thin streams (e.g. Malhan et al., 2018; Ibata et al., 2019, 2021; Li et al., 2022; Ibata et al., 2024) and phase-mixed structures (e.g. Helmi et al., 2018; Belokurov et al., 2018; Haywood et al., 2018; Naidu et al., 2020; Horta et al., 2024). Statistical samples of satellite galaxies (e.g. Mao et al., 2021; Carlsten et al., 2021, 2022; Mao et al., 2024) and bright stellar streams (e.g. Martínez-Delgado et al., 2010, 2023; Miro-Carretero et al., 2024a) have been uncovered around Milky Way-mass galaxies using deep photometric and spectrosopic data, while integrated field spectroscopic surveys have uncovered chemo-dynamical evidence of multiple accretions in inner phase-mixed stellar haloes (Martig et al., 2021; van de Sande et al., 2024).
@@
In addition to encoding the assembly history of the host, disrupting satellites provide unique insights on galaxy formation at small scales and the nature of dark matter. Tidal disruption depends on the internal structural properties of low-mass galaxies, which vary based on dark matter (Du et al., 2018; Tulin & Yu, 2018) and baryonic physics (Sawala et al., 2016; Garrison-Kimmel et al., 2019). Satellites can also disrupt more readily in self-interacting dark matter models due to mass loss from ‘ram pressure stripping’ of dark matter caused by self-interactions with the host (Kaplinghat et al., 2019; Nadler et al., 2020b), resulting in clear differences in the structure of stellar haloes (Forouhar Moreno et al., 2024). The shape and trajectory of stellar streams are extremely sensitive to the local acceleration field (Johnston et al., 2005; Bovy, 2014; Gibbons et al., 2014), making them excellent tracers for measuring the mass and shape of the Milky Way’s dark matter halo and perturbations from massive satellites (Erkal et al., 2019; Shipp et al., 2021; Vasiliev et al., 2021). Such techniques are being extended to external galaxies (Pearson et al., 2022; Nibauer et al., 2023; Walder et al., 2024) in preparation for upcoming deep data from facilities like Rubin (Ivezić et al., 2019), Euclid (Euclid Collaboration et al., 2024), and ARRAKIHS [^1] ARRAKIHS Phase 2 Proposal (public, pdf).
Systems of intact satellites have been extensively studied in cosmological hydrodynamic zoom-in simulations (Brooks & Zolotov, 2014; Wetzel et al., 2016; Sawala et al., 2016; Garrison-Kimmel et al., 2019) and semi-analytic models (Kim et al., 2018; Nadler et al., 2020a) in the context of resolving small-scale challenges to the $\Lambda$CDM framework (Bullock & Boylan-Kolchin, 2017; Sales et al., 2022). In contrast, systematic studies of disrupting satellites around Milky Way-mass haloes are relatively novel. Studies of accreted stellar haloes that identify individual accretion events (Cooper et al., 2010; Deason et al., 2016; Monachesi et al., 2019; Fattahi et al., 2020; Genina et al., 2023) typically do not characterise the level of disruption that each accretion has experienced, instead considering the bound components of accretion events as satellites and all other accreted material as disrupted stellar halo (for an analysis focused on a solar neighbourhood-like volume, see Gómez et al., 2013).
@@ 2 The Auriga Simulations
The original sample of 30 haloes (labelled here as Au-N, with N from 1 to 30) were simulated at ‘level 4’ resolution [^5] The ‘level’ nomenclature traces back to the Aquarius (Springel et al., 2008) dark-matter-only simulations of Milky Way-mass haloes. with baryonic element (dark matter particle) masses of 5.4(29)×$\times 10^{4}$ M and a minimum softening length of 375 pc. Six of these haloes were re-simulated at higher ‘level 3’ resolution (6.7(36)×$\times 10^{3}$ M; 188 pc) and one halo, presented in Grand et al. (2021), at even higher ‘level 2’ resolution (8.5(46)×$\times 10^{2}$ M; 94 pc). Consecutive levels differ by approximately a factor of 8 in mass resolution. We exclude Au-1 and Au-11 from our analysis, as these haloes are experiencing massive mergers at $z=0$. A summary of properties for the Auriga haloes is provided in Table 1.
@@
Level 4 resolution
Au-2 1.915 262 425 8.091 3.072 14.04 8 11 28 0.559 0.026 0.415
Au-3 1.458 239 380 7.477 1.564 9.85 4 8 12 0.199 0.036 0.766
Au-4 1.409 236 393 5.506 3.388 7.13 3 16 17 0.006 0.066 0.928
Au-5 1.186 223 354 6.356 0.876 3.95 4 13 16 0.108 0.169 0.723
Au-6 1.044 214 340 4.796 0.926 7.58 2 5 17 0.310 0.035 0.655
Au-7 1.120 219 358 3.859 2.079 5.05 2 10 14 0.012 0.061 0.927
Au-8 1.081 216 345 3.302 2.252 11.93 5 5 13 0.713 0.191 0.096
Au-9 1.050 214 345 5.875 0.429 3.15 0 10 21 0.000 0.054 0.946
Au-10 1.047 214 338 5.916 0.512 1.90 6 6 15 0.239 0.245 0.516
Au-12 1.093 217 346 5.377 1.484 4.56 2 10 14 0.033 0.334 0.633
Au-13 1.189 223 369 5.830 1.470 2.39 5 12 14 0.377 0.288 0.335
Au-14 1.657 249 392 9.628 2.251 4.58 2 10 16 0.053 0.074 0.873
Au-15 1.222 225 360 3.697 2.194 6.94 6 10 13 0.626 0.118 0.256
Au-16 1.503 241 381 6.290 1.695 12.58 4 12 17 0.526 0.082 0.392
Au-17 1.028 213 336 7.725 0.319 1.91 4 10 21 0.120 0.477 0.403
Au-18 1.221 225 435 8.003 6.091 4.05 12 17 24 0.929 0.018 0.053
Au-19 1.209 225 359 4.786 1.904 7.10 4 8 15 0.255 0.118 0.627
Au-20 1.249 227 385 3.658 3.458 6.10 5 13 16 0.363 0.144 0.494
Au-21 1.451 239 405 6.995 3.557 7.37 8 13 15 0.386 0.281 0.333
Au-22 0.926 205 331 5.999 0.363 1.97 3 5 22 0.421 0.125 0.453
Au-23 1.575 245 387 8.642 1.238 5.19 6 13 24 0.022 0.154 0.825
Au-24 1.492 241 391 6.709 1.477 6.14 7 12 23 0.346 0.251 0.403
Au-25 1.221 225 353 3.433 2.035 9.02 3 9 11 0.017 0.920 0.063
Au-26 1.564 245 406 9.880 2.111 2.17 7 14 21 0.285 0.015 0.699
Au-27 1.745 254 411 8.995 1.843 6.21 4 13 22 0.240 0.098 0.662
Au-28 1.605 247 390 8.465 2.630 3.27 3 12 13 0.015 0.051 0.934
Au-29 1.542 244 410 6.518 5.287 6.38 5 2 23 0.267 0.006 0.727
Au-30 1.108 218 354 3.549 3.412 3.57 0 10 10 0.000 0.690 0.310
Level 3 resolution
Au-6 1.015 212 338 5.801 1.175 4.70 5 11 34 0.321 0.015 0.664
Au-16 1.504 242 381 8.393 2.142 9.44 8 19 45 0.510 0.080 0.410
Au-21 1.415 237 403 7.289 4.716 7.62 10 24 25 0.409 0.283 0.308
Au-23 1.504 242 382 7.896 1.555 7.62 8 19 52 0.017 0.148 0.834
Au-24 1.468 240 389 7.733 1.974 8.25 8 17 52 0.356 0.011 0.633
Au-27 1.696 251 406 8.715 2.298 6.31 5 27 50 0.253 0.179 0.568
Level 2 resolution
Au-6 1.024 212 340 6.495 1.726 6.94 7 11 57 0.335 0.011 0.654
@@
Table 1: Summary properties of Auriga haloes analysed in this work. We provide the halo name; virial mass ($M_{\text{200c}}$) and radius ($R_{\text{200c}}$) defined relative to the critical density; virial radius defined relative to the mean cosmic density ($R_{\text{200m}}$); total in situ ($M_{\ast}^{\text{in situ}}$) and accreted ($M_{\ast}^{\text{acc}}$) stellar mass within $R_{\text{200m}}$ as identified in Section 3.1; spherical radius that encloses half of in situ stellar mass ($R_{50}^{\text{in situ}}$); number of intact satellites ($N_{\text{intact}}$), stellar streams ($N_{\text{stream}}$), and phase-mixed systems ($N_{\text{phase-mixed}}$) as classified in Section 3.2; and the fraction of accreted stellar mass contributed by intact satellites ($f_{\text{acc},\ast}^{\text{intact}}$), stellar streams ($f_{\text{acc},\ast}^{\text{stream}}$), and phase-mixed systems ($f_{\text{acc},\ast}^{\text{phase-mixed}}$).
@@
Halo catalogs are constructed using the on-the-fly SUBFIND halo finder (Springel et al., 2001). This initially identifies haloes using a Friends-of-Friends (FOF) algorithm based on a standard linking length (Davis et al., 1985), then separates them into gravitationally self-bound subhaloes of at least 20 particles. The FOF step is only applied to dark matter particles (other particle types are assigned to the same group as their closest dark matter particle), while SUBFIND is run on all particle types within a FOF group simultaneously. The merger trees were constructed in post-processing with the LHaloTree algorithm (Springel et al., 2005b), which connects FOF and SUBFIND objects both within a single snapshot and across different snapshots. The merger trees are useful for tracking an individual object across time, including when that object merges into another.
@@
It has been shown that the Auriga model produces spiral disc galaxies that are broadly consistent with a number of observations including stellar masses, sizes, and rotation curves (Grand et al., 2017), H i gas distributions (Marinacci et al., 2017), stellar disc warps (Gómez et al., 2017), stellar bars (Fragkoudi et al., 2020, 2024), stellar bulges (Gargiulo et al., 2019), and magnetic fields (Pakmor et al., 2017). In addition, it produces systems of satellite galaxies (Simpson et al., 2018; Grand et al., 2021) and stellar haloes (Monachesi et al., 2019) that are broadly consistent with those observed around Milky Way-mass galaxies and has been used to interpret the Milky Way’s assembly history (Deason et al., 2017; Fattahi et al., 2019), the orbits of its satellites (Riley et al., 2019), and to measure its mass (Callingham et al., 2019; Deason et al., 2019), as well as to produce realistic mock surveys of the Milky Way (Grand et al., 2018; Kizhuprakkat et al., 2024) and M31 (Thomas et al., 2021). The study of Vera-Casanova et al. (2022) is particularly relevant to our work: they characterised the brightest stellar stream around each Auriga host and connected their properties to the overall accreted stellar halo, with a focus on observations from an external perspective.
@@ 3.1 Identifying accretion events
We begin by identifying every star particle within $R_{\text{200m}}$ [^6] This virial radius is defined relative to 200 times the mean cosmic density $\rho_{\mathrm{mean}}=\Omega_{\mathrm{M}}\rho_{\mathrm{crit}}$ and noted as ‘200m.’ ($\sim$350 to $\sim$400 kpc for the Auriga haloes, $\sim$365 kpc for the Milky Way; Deason et al., 2020) of the main host at the present day. This radial extent matches the outer radius considered in Panithanpaisal et al. (2021) and is always larger than the 300 kpc typically considered for observational forecasts of Milky Way satellite galaxies (Newton et al., 2018; Kim et al., 2018; Nadler et al., 2019, 2020a). In addition, Deason et al. (2020) found that the edge of Milky Way-mass galaxies in Auriga and APOSTLE (Sawala et al., 2016; Fattahi et al., 2016), defined as a sharp drop in the logarithmic slope of the stellar density profile, can range from 0.5 to 1×$\times R_{\text{200m}}$, with a typical value of $\sim 0.6R_{\text{200m}}$. This radius adequately captures all accreted stars for our Milky Way-mass haloes, though there is a rich literature on the relationship between the ‘splashback’ radius and a variety of halo properties (Adhikari et al., 2014; Diemer & Kravtsov, 2014; Diemer et al., 2017; Mansfield et al., 2017).

### v2

@@ Abstract
pubyear: 2025pagerange: Auriga Streams I: disrupting satellites surrounding Milky Way-mass haloes at multiple resolutionsC
@@ 1 Introduction
[…]
@@
Figure 1: The accreted stellar mass of the six Auriga galaxies (Au-6, Au-16, Au-21, Au-23, Au-24, Au-27 from left to right) at level 3 resolution, with accretion events split into intact satellites (top), stellar streams (middle), and phase-mixed (bottom) morphologies. The shade corresponds to the amount of stellar mass in a pixel, logarithmically scaled to improve contrast in low/high-density regions. Note that the visible presence of tidal tails from some intact satellites, progenitors for many streams, and cold structure within phase-mixed debris reflects the continuous nature of tidal disruption. The Auriga sample captures a wide range of possible accretion histories for Milky Way-mass haloes, which are efficiently classified by our automated procedure (Section 3.2).
@@
This saga of nearly continuous accretion and disruption of satellites has been observed around both the Milky Way and external galaxies. Wide-area digital sky surveys – in particular the Sloan Digital Sky Survey (York et al., 2000), Dark Energy Survey (DES Collaboration, 2016), and community DECam programs (Drlica-Wagner et al., 2021) – have been incredibly prolific at revealing ever fainter Milky Way satellites (e.g. Willman et al., 2005; Belokurov et al., 2006b; Bechtol et al., 2015; Koposov et al., 2015; Drlica-Wagner et al., 2020; Smith et al., 2024) and stellar streams (e.g. Belokurov et al., 2006a; Koposov et al., 2014; Bernard et al., 2016; Shipp et al., 2018). All-sky astrometry from the Gaia satellite, often in combination with line-of-sight velocities and chemistry from spectrosopic surveys, has produced and characterised an abundance of nearby thin streams (e.g. Malhan et al., 2018; Ibata et al., 2019, 2021; Malhan et al., 2021; Li et al., 2022; Ibata et al., 2024) and phase-mixed structures (e.g. Helmi et al., 2018; Belokurov et al., 2018; Haywood et al., 2018; Naidu et al., 2020; Malhan, 2022; Horta et al., 2024). Statistical samples of satellite galaxies (e.g. Mao et al., 2021; Carlsten et al., 2021, 2022; Mao et al., 2024) and bright stellar streams (e.g. Martínez-Delgado et al., 2010, 2023; Miró-Carretero et al., 2024) have been uncovered around Milky Way-mass galaxies using deep photometric and spectrosopic data, while integrated field spectroscopic surveys have uncovered chemo-dynamical evidence of multiple accretions in inner phase-mixed stellar haloes (Martig et al., 2021; van de Sande et al., 2024).
@@
In addition to encoding the assembly history of the host, disrupting satellites provide unique insights on galaxy formation at small scales and the nature of dark matter. Tidal disruption depends on the internal structural properties of low-mass galaxies, which vary based on dark matter (Du et al., 2018; Tulin & Yu, 2018) and baryonic physics (Sawala et al., 2016; Garrison-Kimmel et al., 2019). Satellites can also disrupt more readily in self-interacting dark matter models due to mass loss from ‘ram pressure stripping’ of dark matter caused by self-interactions with the host (Kaplinghat et al., 2019; Nadler et al., 2020b), resulting in clear differences in the structure of stellar haloes (Moreno et al., 2025). The shape and trajectory of stellar streams are extremely sensitive to the local acceleration field (Johnston et al., 2005; Bovy, 2014; Gibbons et al., 2014; Malhan & Ibata, 2019), making them excellent tracers for measuring the mass and shape of the Milky Way’s dark matter halo and perturbations from massive satellites (Erkal et al., 2019; Shipp et al., 2021; Vasiliev et al., 2021). Such techniques are being extended to external galaxies (Pearson et al., 2022; Nibauer et al., 2023; Walder et al., 2024) in preparation for upcoming deep data from facilities like Rubin (Ivezić et al., 2019), Euclid (Euclid Collaboration et al., 2025), and ARRAKIHS [^1] ARRAKIHS Phase 2 Proposal (public, pdf).
[…]
@@ 2 The Auriga Simulations
[…]
@@
All resolution levels
Au-6-L4 1.044 214 340 4.796 0.926 7.58 2 5 17 0.310 0.035 0.655
Au-6-L3 1.015 212 338 5.801 1.175 4.70 5 11 34 0.321 0.015 0.664
Au-6-L2 1.024 212 340 6.495 1.726 6.94 7 11 57 0.335 0.011 0.654
Levels 4 and 3
Au-16-L4 1.503 241 381 6.290 1.695 12.58 4 12 17 0.526 0.082 0.392
Au-16-L3 1.504 242 381 8.393 2.142 9.44 8 19 45 0.510 0.080 0.410
Au-21-L4 1.451 239 405 6.995 3.557 7.37 8 13 15 0.386 0.281 0.333
Au-21-L3 1.415 237 403 7.289 4.716 7.62 10 24 25 0.409 0.283 0.308
Au-23-L4 1.575 245 387 8.642 1.238 5.19 6 13 24 0.022 0.154 0.825
Au-23-L3 1.504 242 382 7.896 1.555 7.62 8 19 52 0.017 0.148 0.834
Au-24-L4 1.492 241 391 6.709 1.477 6.14 7 12 23 0.346 0.251 0.403
Au-24-L3 1.468 240 389 7.733 1.974 8.25 8 17 52 0.356 0.011 0.633
Au-27-L4 1.745 254 411 8.995 1.843 6.21 4 13 22 0.240 0.098 0.662
Au-27-L3 1.696 251 406 8.715 2.298 6.31 5 27 50 0.253 0.179 0.568
Level 4 only
Au-2-L4 1.915 262 425 8.091 3.072 14.04 8 11 28 0.559 0.026 0.415
Au-3-L4 1.458 239 380 7.477 1.564 9.85 4 8 12 0.199 0.036 0.766
Au-4-L4 1.409 236 393 5.506 3.388 7.13 3 16 17 0.006 0.066 0.928
Au-5-L4 1.186 223 354 6.356 0.876 3.95 4 13 16 0.108 0.169 0.723
Au-7-L4 1.120 219 358 3.859 2.079 5.05 2 10 14 0.012 0.061 0.927
Au-8-L4 1.081 216 345 3.302 2.252 11.93 5 5 13 0.713 0.191 0.096
Au-9-L4 1.050 214 345 5.875 0.429 3.15 0 10 21 0.000 0.054 0.946
Au-10-L4 1.047 214 338 5.916 0.512 1.90 6 6 15 0.239 0.245 0.516
Au-12-L4 1.093 217 346 5.377 1.484 4.56 2 10 14 0.033 0.334 0.633
Au-13-L4 1.189 223 369 5.830 1.470 2.39 5 12 14 0.377 0.288 0.335
Au-14-L4 1.657 249 392 9.628 2.251 4.58 2 10 16 0.053 0.074 0.873
Au-15-L4 1.222 225 360 3.697 2.194 6.94 6 10 13 0.626 0.118 0.256
Au-17-L4 1.028 213 336 7.725 0.319 1.91 4 10 21 0.120 0.477 0.403
Au-18-L4 1.221 225 435 8.003 6.091 4.05 12 17 24 0.929 0.018 0.053
Au-19-L4 1.209 225 359 4.786 1.904 7.10 4 8 15 0.255 0.118 0.627
Au-20-L4 1.249 227 385 3.658 3.458 6.10 5 13 16 0.363 0.144 0.494
Au-22-L4 0.926 205 331 5.999 0.363 1.97 3 5 22 0.421 0.125 0.453
Au-25-L4 1.221 225 353 3.433 2.035 9.02 3 9 11 0.017 0.920 0.063
Au-26-L4 1.564 245 406 9.880 2.111 2.17 7 14 21 0.285 0.015 0.699
Au-28-L4 1.605 247 390 8.465 2.630 3.27 3 12 13 0.015 0.051 0.934
Au-29-L4 1.542 244 410 6.518 5.287 6.38 5 2 23 0.267 0.006 0.727
Au-30-L4 1.108 218 354 3.549 3.412 3.57 0 10 10 0.000 0.690 0.310
@@
Table 1: Summary properties of Auriga haloes analysed in this work. We provide the halo name and resolution level; virial mass ($M_{\text{200c}}$) and radius ($R_{\text{200c}}$) defined relative to the critical density; virial radius defined relative to the mean cosmic density ($R_{\text{200m}}$); total in situ ($M_{\ast}^{\text{in situ}}$) and accreted ($M_{\ast}^{\text{acc}}$) stellar mass within $R_{\text{200m}}$ as identified in Section 3.1; spherical radius that encloses half of in situ stellar mass ($R_{50}^{\text{in situ}}$); number of intact satellites ($N_{\text{intact}}$), stellar streams ($N_{\text{stream}}$), and phase-mixed systems ($N_{\text{phase-mixed}}$) as classified in Section 3.2; and the fraction of accreted stellar mass contributed by intact satellites ($f_{\text{acc},\ast}^{\text{intact}}$), stellar streams ($f_{\text{acc},\ast}^{\text{stream}}$), and phase-mixed systems ($f_{\text{acc},\ast}^{\text{phase-mixed}}$).
@@
Halo catalogs are constructed using the on-the-fly SUBFIND halo finder (Springel et al., 2001). This initially identifies haloes using a Friends-of-Friends (FOF) algorithm based on a standard linking length (Davis et al., 1985), then separates them into gravitationally self-bound subhaloes of at least 20 particles. The FOF step is only applied to dark matter particles (other particle types are assigned to the same group as their closest dark matter particle), while SUBFIND is run on all particle types within a FOF group simultaneously. The merger trees were constructed in post-processing with the LHaloTree algorithm (Springel et al., 2005b), which connects FOF and SUBFIND objects both within a single snapshot and across different snapshots. The merger trees are useful for tracking an individual object across time, including when that object merges into another. LHaloTree identifies the descendant of a target halo by tracking which halo contains the most dark matter particles from the target in the preceding snapshot, giving higher weight to particles that were more tightly bound to the target.
@@
It has been shown that the Auriga model produces spiral disc galaxies that are broadly consistent with a number of observations including stellar masses, sizes, and rotation curves (Grand et al., 2017), H i gas distributions (Marinacci et al., 2017), stellar disc warps (Gómez et al., 2017), stellar bars (Fragkoudi et al., 2020, 2025), stellar bulges (Gargiulo et al., 2019), and magnetic fields (Pakmor et al., 2017). In addition, it produces systems of satellite galaxies (Simpson et al., 2018; Grand et al., 2021) and stellar haloes (Monachesi et al., 2019) that are broadly consistent with those observed around Milky Way-mass galaxies and has been used to interpret the Milky Way’s assembly history (Deason et al., 2017; Fattahi et al., 2019), the orbits of its satellites (Riley et al., 2019), and to measure its mass (Callingham et al., 2019; Deason et al., 2019), as well as to produce realistic mock surveys of the Milky Way (Grand et al., 2018; Kizhuprakkat et al., 2024) and M31 (Thomas et al., 2021). The study of Vera-Casanova et al. (2022) is particularly relevant to our work: they characterised the brightest stellar stream around each Auriga host and connected their properties to the overall accreted stellar halo, with a focus on observations from an external perspective.
@@ 3.1 Identifying accretion events
[…]